The Chance of Dying of a disease is 3.5% without a drug and 7% with the...

  1. The Chance of Dying of a disease is 3.5% without a drug and 7% with the drug. Compute the risks, odds, and their respective ratios and write out in words what each means?
Answer #1

Percentage of patients with chances of dying of a disease without a drug - 3.5%

Percentage of patients with chances of dying of disease with a drug - 7%

Total group at risk = 10.5 (7+3.5)

Relative risk for patients with chances of getting cured without drugs = 3.5/10.5 = 0.333

So for 100 percent of patients with disease 33.3 percent lpatients have the chances to get cured without drugs.

Relative risk for patients with chances of getting cured with drugs = 7.5/10.5 = 0.714

And for patients with drugs they have a chances of 71.4 percent of getting cured

Odds ratio for this risk 3.5/10.5/7/10.5 = 0.5

This suggest that people are more probably to get cured by drugs and less likely without drugs.